OpenStudy (anonymous):

What is the difference? Express your answer in lowest terms. -7/15 - (-11/15)

OpenStudy (anonymous):

@Susanna @mathmale

OpenStudy (anonymous):

ok i got 82/255

OpenStudy (susanna):

\[-\frac{ 7 }{ 15 } + \frac{ 11 }{ 15 }\]

OpenStudy (anonymous):

wait what

OpenStudy (anonymous):

i thought it was minusing

OpenStudy (susanna):

\[-\frac{ 6 }{ 15 }\]

OpenStudy (susanna):

It was, but you subtracted a negative number, and two negatives make a positive. for example, 5-(-2)=5+2=7

OpenStudy (anonymous):

ohh ok i for got that hxs now simplifiy?

OpenStudy (susanna):

yes simplyfy -6/15

OpenStudy (anonymous):

ok 2/6 2/15?

OpenStudy (susanna):

Divide both the numerator and denominator by the same number.

OpenStudy (susanna):

\[-\frac{ 6 }{ 15 }=-\frac{ 2 }{ 5 }\]

OpenStudy (susanna):

divided by 3

OpenStudy (anonymous):

ok thxs i have more

OpenStudy (susanna):

sure